A long straight wire of radius R carries a steady current I. The current is uniformly distributed over its cross section. The magnetic field at a distance 2a/3 from its centre is


1
Expert's answer
2021-07-06T11:40:02-0400

We know that

"\oint B.dl=\mu_0i"

"B\oint dl=\mu_0i=B.(2\pi a)=\mu_0i"

"B=\frac{\mu_0i}{2\pi a}"

Magnetic field at centre of wire

"B=\frac{\mu_0i}{2\pi a}"

When

"r=\frac{2a}{3}"

"\oint B'.dl=\mu_0i_{en}"

"i_{en}=i\frac{\pi \times( \frac{2a}{3})^2}{\pi a^2}=\frac{4}{9}i"

"B'(2\pi\times\frac{2a}{3})=\mu_0\frac{4i}{9}"

"B'=\frac{\mu_0 I}{3 \pi a}"

"\frac{B'}{B}=\frac{2}{3}"

"B'=\frac{2}{3}B"